OpenStudy (anonymous):

I need help with this problem: Bayside Insurance offers two health plans. Under plan A, Giselle would have to pay the first $160 of her medical bills, plus 25% of the rest. Under plan B, Giselle would pay the first $230, but only 20% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $230 in bills. Giselle would save with plan B if she had more than how much in bills

OpenStudy (kropot72):

Let r = the rest of the bill The break-even point for the plans will arise when: 160 + 0.25r = 230 + 0.2r 0.25r = (230-160) + 0.2r 0.05r = 70 r = 1400 Therefore the bill = 160 + 1400 = $1560. Plan B will save money for a bill greater than $1560

OpenStudy (anonymous):

@ kropot72 Its saying the the answer is not correct

OpenStudy (kropot72):

I am rechecking. Please wait :)

OpenStudy (pfenn1):

is 1280 the answer?

OpenStudy (kropot72):

The correct answer is for bills greater than $1280 plan A will save. I will post the solution. Please wait.

OpenStudy (anonymous):

Yes @ pfenni the anwer is 1280 here is another one: Bayside Insurance offers two health plans. Under plan A, Giselle would have to pay the first $50 of her medical bills, plus 30% of the rest. Under plan B, Giselle would pay the first $110, but only 20% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $110 in bills. Giselle would save with plan B if she had more than how much in bills

OpenStudy (kropot72):

Let A = amount of bill. Then for the plans to break even: 160 + 0.25(A-160) = 230 + 0.2(A-230) 0.05A = 64 A = 1280

OpenStudy (anonymous):

here is another one: Bayside Insurance offers two health plans. Under plan A, Giselle would have to pay the first $50 of her medical bills, plus 30% of the rest. Under plan B, Giselle would pay the first $110, but only 20% of the rest. For what amount of medical bills will plan B save Giselle money? Assume she has over $110 in bills. Giselle would save with plan B if she had more than how much in bills

OpenStudy (kropot72):

For the second problem the break-even point for the plans is when: 50 + 0.3(A-50) = 110 + 0.2(A-110) From which we find that plan B will save for bills more than $530

OpenStudy (anonymous):

Thanks!

OpenStudy (kropot72):

You're welcome :)